When light with wavelength 740 nm shines on a particular sheet of metal, it ejects electrons that have a maximum kinetic energy of 0.4 eV . What is the maximum kinetic energy of ejected electrons if we shine light with wavelength 380 nm  on the sheet of metal instead?
Answer #1

t.= 740x10ºm, ki z ouer, to = 380x10-m, kz=? Since E = K+P = h = K+Q & Q = he look The kat (hele ki) Ka - he + KA 3) Kx = 12